# Algorithms for mean-field variational inference via polyhedral optimization in the Wasserstein space

### v1

@@ Abstract
We develop a theory of finite-dimensional polyhedral subsets over the Wasserstein space and optimization of functionals over them via first-order methods. Our main application is to the problem of mean-field variational inference, which seeks to approximate a distribution $\pi$ over $\mathbb{R}^{d}$ by a product measure $\pi^{\star}$. When $\pi$ is strongly log-concave and log-smooth, we provide (1) approximation rates certifying that $\pi^{\star}$ is close to the minimizer $\pi^{\star}_{\diamond}$ of the KL divergence over a polyhedral set $\mathcal{P}_{\diamond}$, and (2) an algorithm for minimizing $\operatorname*{KL}(\cdot\!\;\|\;\!\pi)$ over $\mathcal{P}_{\diamond}$ with accelerated complexity $O(\sqrt{\kappa}\log(\kappa d/\varepsilon^{2}))$, where $\kappa$ is the condition number of $\pi$.
@@ 1 Introduction
Optimization over (subsets of) the Wasserstein space (the metric space of probability measures over $\mathbb{R}^{d}$ endowed with the $2$-Wasserstein distance, see Section 2) has found diverse and effective applications in modern machine learning. Notable examples include distributionally robust optimization (Kuhn et al.,, 2019; Yue et al.,, 2022), the computation of barycenters (Cuturi and Doucet,, 2014; Zemel and Panaretos,, 2019; Chewi et al.,, 2020; Altschuler et al.,, 2021; Backhoff-Veraguas et al.,, 2022), sampling (Jordan et al.,, 1998; Wibisono,, 2018; Chewi,, 2023), and variational inference (see below). The development of optimization algorithms over this space, however, has been hindered by significant implementation challenges stemming from its infinite-dimensional nature and the curse of dimensionality which impedes efficient representation of high-dimensional distributions.
@@
In this paper, we instead study the problem of mean-field VI, in which $\mathcal{C}$ is taken to be the class of product measures over $\mathbb{R}^{d}$, written $\mathcal{P}(\mathbb{R})^{\otimes d}$. In this context, the works Yao and Yang, (2022); Lacker, (2023) have also developed algorithms based on Wasserstein gradient flows, although computational guarantees for VI are still nascent (see Section 5.1 for further details and comparison with the literature).
@@
We discuss the implementation of our algorithm in Section 5.5.1. Below, we describe our contributions in more detail.
@@
We next turn toward mean-field VI as a compelling application of our theory of polyhedral optimization. Throughout, we only assume that $\pi$ satisfies the standard assumptions of strong log-concavity and log-smoothness. By leveraging the structure of the mean-field VI solution and establishing regularity bounds for optimal transport maps between well-conditioned product measures, we first prove an approximation result which shows that the solution $\pi^{\star}$ to mean-field VI in (1) is well-approximated by the minimizer $\pi_{\diamond}^{\star}$ of the KL divergence over of a suitable polyhedral approximation $\mathcal{P}_{\diamond}$ of the space of product measures. Importantly, our approximation rates, owing to the coordinate-wise decomposability of mean-field VI, do not incur the curse of dimensionality.
Next, we establish the geodesic strong convexity and geodesic smoothness of the KL divergence over $\mathcal{P}_{\diamond}$. Consequently, bringing to bear the full force of the Euclidean–Wasserstein equivalence, we obtain, to the best of our knowledge, the first accelerated and end-to-end convergence rates for mean-field VI.
@@
Our paper is outlined as follows: In Section 2 we provide relevant background information on optimal transport theory and the Wasserstein space. In Section 3, we develop known properties of compatibility which provide tools for building large compatibile families, and we establish the key isometry with (+$(\mathbb{R}^{|\mathcal{M}|}_{+},\|\cdot\|_{Q})$. We prove our general continuous- and discrete-time guarantees for Wasserstein polyhedral optimization in Section 4. We apply our framework to mean-field variational inference in Section 5; in particular, the implementation is discussed in Section 5.5.1. Finally, we conclude with numerous open directions.
@@
where $\Delta_{|\mathcal{M}|}$ is the $|\mathcal{M}|$-simplex. Such a constraint set is used in Boissard et al., (2015); Gunsilius et al., (2022); Werenski et al., (2022), and is usually studied in the context of Wasserstein barycenters. The work of Bonneel et al., (2016) also considers $\operatorname{conv}(\mathcal{M})_{\sharp}\rho$, but makes no assumptions on the maps, and they tackle the problem from a computational angle via Sinkhorn’s algorithm (Cuturi,, 2013; Peyré and Cuturi,, 2019), albeit without convergence guarantees. Albergo et al., (2023) use the same set, but without incorporating any optimal transport theory.
@@ 3.2 Isometry with Euclidean geometry
where we applied compatibility in the last equality to argue that $\tilde{T}\circ T^{-1}$ is the optimal transport map from $T_{\sharp}\rho$ to $\tilde{T}_{\sharp}\rho$. This equality shows that for compatible $\mathcal{M}$, the geometry of $\operatorname{cone}(\mathcal{M})_{\sharp}\rho$ is in a sense trivial, being isometric to a convex subset of the Hilbert space $L^{2}(\rho)$. This fundamental property lies at the heart of the widespread usage of one-dimensional optimal transport in applications, see Cai et al., (2020).
@@
Let $\mathcal{M}$ be a finite family of optimal transport maps. Then, (+$(\mathbb{R}^{|\mathcal{M}|}_{+},\|\cdot\|_{Q})$ is always isometric to $(\operatorname{cone}(\mathcal{M})_{\sharp}\rho,\mathsf{d}_{\mathrm{LOT},\rho})$. If, in addition, $\mathcal{M}$ is a compatible family (i.e., $\operatorname{cone}(\mathcal{M})_{\sharp}\rho$ is polyhedral), then (+$(\mathbb{R}^{|\mathcal{M}|}_{+},\|\cdot\|_{Q})$ is isometric to $(\operatorname{cone}(\mathcal{M})_{\sharp}\rho,W_{2})$.
@@ 4.1 Continuous-time gradient flow
The isometry of Section 3.2 implies that the constrained Wasserstein gradient flow of $\mathcal{F}$ is equivalent to the gradient flow of the functional $\lambda\mapsto\mathcal{F}(\mu_{\lambda})$ with respect to the $Q$-geometry. The latter gradient flow can be written explicitly as
@@
Then, geodesic strong convexity over $\mathbb{W}$ translates to strong convexity of $\lambda\mapsto\mathcal{F}(\mu_{\lambda})$ over $(\mathbb{R}_{+}^{|\mathcal{M}|},\|\cdot\|_{Q})$ for free. The following theorem establishes convergence rates for this continuous-time flow; its proof is found in Lambert et al., (2022, Appendix D).
@@ 4.2.1 Accelerated projected gradient descent
Let $\operatorname{cone}(\mathcal{M})_{\sharp}\rho$ be polyhedral and $\mathcal{K}\subseteq\operatorname{cone}(\mathcal{M})$ be convex. Suppose that $\mathcal{F}$ is $m$-strongly convex and $M$-smooth over $\mathcal{K}_{\sharp}\rho$ and let $\mu_{\star}$ denote a minimizer over this set. Let $(\lambda^{(t)}:t=0,1,2,3\ldots)$ denote the iterates of Algorithm 1.
@@ 1. $(\mathsf{VB})$
    $\mathbb{E}[\|Q^{-1}\,(\hat{\nabla}_{\lambda}{\mathcal{F}}(\mu_{\lambda})-\nabla_{\lambda}{\mathcal{F}}(\mu_{\lambda}))\|_{Q}^{2}]\leq c_{0}+c_{1}\,\mathbb{E}W_{2}^{2}(\mu_{\lambda},\mu_{\star})\,.$
@@
Let $\operatorname{cone}(\mathcal{M})_{\sharp}\rho$ be polyhedral and $\mathcal{K}\subseteq\operatorname{cone}(\mathcal{M})$ be convex. Suppose that $\mathcal{F}$ is $m$-strongly convex and $M$-smooth over $\mathcal{K}_{\sharp}\rho$, let $\mu_{\star}$ denote a minimizer over this set, and suppose that $(\mathsf{VB})$ holds. Let $(\lambda^{(t)}:t=0,1,2,3\ldots)$ denote the iterates of Algorithm 2. If we choose $h\asymp\frac{m\varepsilon^{2}}{c_{0}}\leq\frac{m}{2c_{1}}\wedge\frac{1}{2\kappa M}$, and the number of iterations is at least
@@
then $\mathbb{E}W_{2}^{2}(\mu_{\lambda^{(t)}},\mu_{\star})\leq\varepsilon^{2}$.
@@ 4.2.3 Frank–Wolfe
Suppose that $\mathcal{F}$ is geodesically convex and $M$-smooth over $\operatorname{conv}(\mathcal{M})_{\sharp}\rho$, and let $\mu_{\star}$ be a minimizer of $\mathcal{F}$ over this set. Let $(\lambda^{(t)}:t=0,1,2,3\ldots)$ denote the iterates of Algorithm 3, with Choose $\alpha^{(t)}=2/(t+2)$. Then,
@@
     $j^{*}\leftarrow\operatorname*{arg\,min}_{j\in[J]}{\langle\nabla_{\lambda}\mathcal{F}(\mu_{\lambda^{(t)}}),e_{j}-\lambda^{(t)}_{j}\rangle}$
@@ 4.3 Enriching the family of compatible maps
The augmented cone is parameterized by $(\lambda,v)\in\mathbb{R}_{+}^{|\mathcal{M}|}\times\mathbb{R}^{d}$. We may assume that each of the maps $T\in\mathcal{M}$ has mean zero under $\rho$, since this does not affect the augmented cone. Under this assumption, it is easy to see (c.f. the proof of 5.10) that we still obtain an isometry with a Euclidean metric: $W_{2}^{2}(\mu_{\eta,u},\mu_{\lambda,v})=\|\eta-\lambda\|_{Q}^{2}+\|u-v\|^{2}$. In this setting, the first-order algorithms must be modified to compute the gradient and projection steps with respect to this metric.
@@ 5 Application to mean-field variational inference
where $Z$, the unknown normalizing constant of $\pi\propto\exp(-V)$ plays no part in the optimization problem. The following assumption, which will play a crucial role in our analyses in Section 5.3 and Section 5.4, is standard in the literature on log-concave sampling (Chewi,, 2023):
@@ 5.1 Mean-field variational inference
Mean-field VI has a rich history in the realm of statistical inference; see Section 2.3 in Blei et al., (2017) for a brief historical introduction. Despite being widely used, computational and statistical guarantees have only recently emerged. A standard algorithm to solve (12) is Coordinate Ascent VI (CAVI) (see Blei et al.,, 2017, Section 2.4), the updates for which can be implemented easily in certain conjugate models. Guarantees for CAVI were provided recently in Bhattacharya et al., (2023) under a generalized correlation condition for $\pi$.
More closely related to our work is the use of Wasserstein gradient flows. The work of Lacker, (2023) connects mean-field VI to constrained Wasserstein gradient flows, providing continuous-time guarantees via projected log-Sobolev inequalities but without a concrete algorithmic implementation. Also, in the context of a Bayesian latent variable models, convergence guarantees for a Wasserstein gradient flow under a well-conditioned assumption at the population level was established by Yao and Yang, (2022). Toward the issue of implementation, they suggested two strategies based on particle approximation combined with either Langevin sampling or optimization over transport maps respectively, but they did not analyze the error arising from the particle approximation. Despite the promising nature of these works, implementation remains a challenge. [^3] Tran et al., (2023) also study discrete-time rates for mean-field VI, though their draft has numerous errors, and incorrectly applies several prior results.
@@ 5.2 Regularity of optimal transport maps between well-conditioned product measures
[…]

### v4

@@ Abstract
We develop a theory of finite-dimensional polyhedral subsets over the Wasserstein space and optimization of functionals over them via first-order methods. Our main application is to the problem of mean-field variational inference (MFVI), which seeks to approximate a distribution $\pi$ over $\mathbb{R}^{d}$ by a product measure $\pi^{\star}$. When $\pi$ is strongly log-concave and log-smooth, we provide (1) approximation rates certifying that $\pi^{\star}$ is close to the minimizer $\pi^{\star}_{\diamond}$ of the KL divergence over a polyhedral set $\mathcal{P}_{\diamond}$, and (2) an algorithm for minimizing $\operatorname*{KL}(\cdot\!\;\|\;\!\pi)$ over $\mathcal{P}_{\diamond}$ based on accelerated gradient descent over $\mathbb{R}^{d}$. As a byproduct of our analysis, we obtain the first end-to-end analysis for gradient-based algorithms for MFVI.
@@ 1 Introduction
Optimization over (subsets of) the Wasserstein space (the metric space of probability measures over $\mathbb{R}^{d}$ endowed with the $2$-Wasserstein distance, see Section 2) has found diverse and effective applications in modern machine learning. Notable examples include distributionally robust optimization (Kuhn et al.,, 2019; Yue et al.,, 2022), the computation of barycenters (Cuturi and Doucet,, 2014; Zemel and Panaretos,, 2019; Chewi et al.,, 2020; Altschuler et al.,, 2021; Backhoff-Veraguas et al.,, 2022), sampling (Jordan et al.,, 1998; Wibisono,, 2018; Chewi,, 2024), and variational inference (see below). The development of optimization algorithms over this space, however, has been hindered by significant implementation challenges stemming from its infinite-dimensional nature and the curse of dimensionality which impedes efficient representation of high-dimensional distributions.
@@
In this paper, we instead study the problem of mean-field VI, in which $\mathcal{C}$ is taken to be the class of product measures over $\mathbb{R}^{d}$, written $\mathcal{P}(\mathbb{R})^{\otimes d}$. In this context, the works Zhang and Zhou, (2020); Yao and Yang, (2022); Lacker, (2023) have also developed algorithms based on Wasserstein gradient flows, although computational guarantees for VI are still nascent (see Section 5.1 for further details and comparison with the literature).
@@
We discuss the implementation of our algorithm in Section 5.5.1, with code available here. Below, we describe our contributions in more detail.
@@
We next turn toward mean-field VI as a compelling application of our theory of polyhedral optimization. Throughout, we only assume that $\pi$ satisfies the standard assumptions of strong log-concavity and log-smoothness. By leveraging the structure of the mean-field VI solution and establishing regularity bounds for optimal transport maps between well-conditioned product measures, we first prove an approximation result which shows that the solution $\pi^{\star}$ to mean-field VI in (1) is well-approximated by the minimizer $\pi_{\diamond}^{\star}$ of the KL divergence over a suitable polyhedral approximation $\mathcal{P}_{\diamond}$ of the space of product measures. Importantly, our approximation rates, owing to the coordinate-wise decomposability of mean-field VI, do not incur the curse of dimensionality.
Next, we establish the geodesic strong convexity and geodesic smoothness of the KL divergence over $\mathcal{P}_{\diamond}$. Consequently, bringing to bear the full force of the Euclidean–Wasserstein equivalence, we obtain, to the best of our knowledge, the first end-to-end convergence rates for mean-field VI.
@@
Our paper is outlined as follows: In Section 2 we provide relevant background information on optimal transport theory and the Wasserstein space. In Section 3, we develop known properties of compatibility which provide tools for building large compatible families, and we establish the key isometry with (+$(\mathbb{R}^{|\mathcal{M}|}_{+},\|\cdot\|_{Q})$. We prove our general continuous- and discrete-time guarantees for Wasserstein polyhedral optimization in Section 4. We apply our framework to mean-field variational inference in Section 5; in particular, the implementation is discussed in Section 5.5.1. Section 6 contains our numerical experiments. Section 7 extends our framework to mixtures of product measures. We conclude with numerous open directions.
@@
where $\Delta_{|\mathcal{M}|}$ is the $|\mathcal{M}|$-simplex. Such a constraint set is used by Boissard et al., (2015); Gunsilius et al., (2022); Werenski et al., (2022), and is usually studied in the context of Wasserstein barycenters. The work of Bonneel et al., (2016) also considers $\operatorname{conv}(\mathcal{M})_{\sharp}\rho$, but makes no assumptions on the maps, and they tackle the problem from a computational angle via Sinkhorn’s algorithm (Cuturi,, 2013; Peyré and Cuturi,, 2019), albeit without convergence guarantees. Albergo et al., (2024) use the same set, but without incorporating any optimal transport theory.
@@
Finally, we mention that our work falls under the category of linearized optimal transport (Wang et al.,, 2013), which we closely address in Section 3.2.
@@ 3.2 Isometry with Euclidean geometry
where we applied compatibility in the last equality to argue that $\tilde{T}\circ T^{-1}$ is the optimal transport map from $T_{\sharp}\rho$ to $\tilde{T}_{\sharp}\rho$. This equality shows that for compatible $\mathcal{M}$, the geometry of $\operatorname{cone}(\mathcal{M})_{\sharp}\rho$ is in a sense trivial, being isometric to a convex subset of the Hilbert space $L^{2}(\rho)$. This fundamental property lies at the heart of the widespread usage of one-dimensional optimal transport in applications, see Wang et al., (2013); Basu et al., (2014); Kolouri and Rohde, (2015); Kolouri et al., (2016); Park and Thorpe, (2018); Cai et al., (2020); Khurana et al., (2023) for applications.
@@
Let $\mathcal{M}$ be a finite family of optimal transport maps with $Q$ defined as the Gram matrix with entries $Q_{T,\tilde{T}}=\langle T,\tilde{T}\rangle_{L^{2}(\rho)}$ for $T,\tilde{T}\in\mathcal{M}$. Then, (+$(\mathbb{R}^{|\mathcal{M}|}_{+},\|\cdot\|_{Q})$ is always isometric to $(\operatorname{cone}(\mathcal{M})_{\sharp}\rho,\mathsf{d}_{\mathrm{LOT},\rho})$. If, in addition, $\mathcal{M}$ is a compatible family (i.e., $\operatorname{cone}(\mathcal{M})_{\sharp}\rho$ is polyhedral), then (+$(\mathbb{R}^{|\mathcal{M}|}_{+},\|\cdot\|_{Q})$ is isometric to $(\operatorname{cone}(\mathcal{M})_{\sharp}\rho,W_{2})$.
@@ 4.1 Continuous-time gradient flow
The isometry of Section 3.2 implies that the constrained Wasserstein gradient flow of $\mathcal{F}$ is equivalent to the gradient flow of the functional $\lambda\mapsto\mathcal{F}(\mu_{\lambda})$ with respect to the $Q$-geometry. [^2] See Nesterov, (2018, §4.2.1) for a thorough discussion on optimization over general Euclidean spaces. The latter gradient flow can be written explicitly as
@@
Then, geodesic strong convexity over $\mathbb{W}$ translates to strong convexity of $\lambda\mapsto\mathcal{F}(\mu_{\lambda})$ over $(\mathbb{R}_{+}^{|\mathcal{M}|},\|\cdot\|_{Q})$ for free. The following theorem [^3] In the case where we further constrain the gradient flow to lie in a convex set, (7) should be replaced by a differential inclusion. Since this is not relevant to the subsequent developments, we omit a fuller discussion of this point. establishes convergence rates for this continuous-time flow; see Lambert et al., (2022, Appendix D) for a proof.
@@ 4.2.1 Accelerated projected gradient descent
Let $\operatorname{cone}(\mathcal{M})_{\sharp}\rho$ be polyhedral and $\mathcal{K}\subseteq\operatorname{cone}(\mathcal{M})$ be convex. Suppose that $\mathcal{F}$ is geodesically $m$-strongly convex and $M$-smooth over $\mathcal{K}_{\sharp}\rho$ and let $\mu_{\star}$ denote a minimizer over this set. Let $(\lambda^{(t)}:t=0,1,2,3\ldots)$ denote the iterates of Algorithm 1.
@@ 1. $(\mathsf{VB})$
    $\mathbb{E}[\|Q^{-1}\,(\hat{\nabla}_{\lambda}{\mathcal{F}}(\mu_{\lambda})-\nabla_{\lambda}{\mathcal{F}}(\mu_{\lambda}))\|_{Q}^{2}]\leq c_{0}+c_{1}\,\mathbb{E}[W_{2}^{2}(\mu_{\lambda},\mu_{\star})]\,.$
@@
Let $\operatorname{cone}(\mathcal{M})_{\sharp}\rho$ be polyhedral and $\mathcal{K}\subseteq\operatorname{cone}(\mathcal{M})$ be convex. Suppose that $\mathcal{F}$ is geodesically $m$-strongly convex and $M$-smooth over $\mathcal{K}_{\sharp}\rho$, let $\mu_{\star}$ denote a minimizer over this set, and suppose that $(\mathsf{VB})$ holds. Let $(\lambda^{(t)}:t=0,1,2,3\ldots)$ denote the iterates of Algorithm 2 and let $\varepsilon>0$ be sufficiently small. If we choose step size $h\asymp\frac{m\varepsilon^{2}}{c_{0}}\leq\frac{m}{2c_{1}}\wedge\frac{1}{2\kappa M}$, and the number of iterations is at least
@@
then $\mathbb{E}[W_{2}^{2}(\mu_{\lambda^{(t)}},\mu_{\star})]\leq\varepsilon^{2}$.
@@ 4.2.3 Frank–Wolfe
Suppose that $\mathcal{F}$ is geodesically convex and $M$-smooth over $\operatorname{conv}(\mathcal{M})_{\sharp}\rho$, and let $\mu_{\star}$ be a minimizer of $\mathcal{F}$ over this set. Let $(\lambda^{(t)}:t=0,1,2,3\ldots)$ denote the iterates of Algorithm 3, with step size $\alpha^{(t)}=2/(t+2)$. Then,
@@
     $j^{*}\leftarrow\operatorname*{arg\,min}_{j\in[J]}{\langle\nabla_{\lambda}\mathcal{F}(\mu_{\lambda^{(t)}}),e_{j}-\lambda^{(t)}\rangle}$
@@ 4.3 Enriching the family of compatible maps
The augmented cone is parameterized by $(\lambda,v)\in\mathbb{R}_{+}^{|\mathcal{M}|}\times\mathbb{R}^{d}$. We may assume that each of the maps $T\in\mathcal{M}$ has mean zero under $\rho$, since this does not affect the augmented cone. Under this assumption, it is easy to see (c.f. the proof of 5.11) that we still obtain an isometry with a Euclidean metric: $W_{2}^{2}(\mu_{\eta,u},\mu_{\lambda,v})=\|\eta-\lambda\|_{Q}^{2}+\|u-v\|^{2}$. In this setting, the first-order algorithms must be modified to compute the gradient and projection steps with respect to this metric.
Remark 4.6 (Broader impact of our framework).
We now pause to briefly discuss the broader impact of polyhedral sets. We want to stress that, even without compatibility, our framework can be used to optimize functionals over any convex subset of the tangent space, provided that the functional is convex with respect to the linearized optimal transport distance. In turn, this is equivalent to requiring that the functional is convex along generalized geodesics (see Ambrosio et al.,, 2005, §9.2), which is typically the case when the functional is convex in the Wasserstein geometry; for example, it holds for the KL divergence with respect to a log-concave measure. This substantially expands the scope of applications as it allows for optimization over any convex subset of the tangent space, not just compatible ones.
@@ 5 Application to mean-field variational inference
where $Z$, the unknown normalizing constant of $\pi\propto\exp(-V)$ plays no part in the optimization problem. The following assumption, which will play a crucial role in our analyses in Section 5.3 and Section 5.4, is standard in the literature on log-concave sampling (Chewi,, 2024):
@@ 5.1 Mean-field variational inference
Mean-field VI has a rich history in the realm of statistical inference; see Section 2.3 in Blei et al., (2017) for a brief historical introduction. Despite being widely used, computational and statistical guarantees have only recently emerged. A standard algorithm to solve (12) is Coordinate Ascent VI (CAVI) (see Blei et al.,, 2017, Section 2.4), the updates for which can be implemented for certain conjugate models. Guarantees for CAVI were provided recently in Bhattacharya et al., (2023) under a generalized correlation condition for $\pi$; see also Arnese and Lacker, (2024) and Lavenant and Zanella, (2024).
More closely related to our work is the use of Wasserstein gradient flows. The work of Lacker, (2023) connects mean-field VI to constrained Wasserstein gradient flows, providing continuous-time guarantees via projected log-Sobolev inequalities but without a concrete algorithmic implementation; see also Lacker et al., (2024). Also, in the context of a Bayesian latent variable models, convergence guarantees for a Wasserstein gradient flow under a well-conditioned assumption at the population level was established by Yao and Yang, (2022). Toward the issue of implementation, they suggested two strategies based on particle approximation combined with either Langevin sampling or optimization over transport maps respectively, but they did not analyze the error arising from the particle approximation. Zhang and Zhou, (2020) study the theoretical and computational properties of mean-field variational inference in the context of community detection. Despite the promising nature of these works, implementation remains a challenge in complete generality.
@@ 5.2 Regularity of optimal transport maps between well-conditioned product measures
$|(T^{\star}_{i})^{\prime\prime}(x)|$ $\lesssim\frac{\kappa}{\sqrt{\ell_{V}}}\,(1+|x|)\,,\qquad\text{and}\qquad|(T^{\star}_{i})^{\prime\prime\prime}(x)|\lesssim\frac{\kappa^{2}}{\sqrt{\ell_{V}}}\,(1+|x|^{2})\,.$ (16)